dimensions of A and B when given the formula speed is equals to A multiplied by acceleration due to gravity squared times density to the power negative one plus B squared multiplied by t 



1
Expert's answer
2021-11-01T12:54:00-0400

"v=Ag^2\rho^{-1}+B^2t,"

"[A]={v\rho }{g^{-2}}=1~\frac ms\cdot \frac{kg}{m^3}\cdot \frac{s^4}{m^2}=1~\frac{kg\cdot s^3}{m^4}."

"[B]=\sqrt{vt^{-1}}=1~\frac{\sqrt m}s."